# Higher Memory Effects in Numerical Simulations of Binary Black Hole Mergers

### v2

@@ V Discussion
In this work we provide straightforward expressions for the first three moments of the news (that is, the displacement, drift, and first higher memories) in terms of charges and radiative and non-radiative fluxes as well as the first numerical calculation of these terms in the context of a binary black hole merger. Besides providing a hierarchy for these terms in terms of how much they contribute to the various memory effects, we also explore how much the morphology of these contributions match “derivatives of step functions”, which is what they should resemble if one assumes that they predominantly source non-oscillatory memory effects. In particular, we find that while the $\mathcal{F}_{0}$ and $\mathcal{F}_{1}$ flux contributions to the displacement and spin memory effects resemble step- and delta-like functions, the other fluxes break from this expected behavior. A possible explanation for this, however, could simply be that to make these terms exhibit this behavior, then one must perform a correction like the Wald-Zoupas correction that is used for the first moment of the news Wald and Zoupas (2000); Flanagan and Nichols (2017); Grant et al. (2022) or remove the nontrivial QNM contributions that are sourced during the ringdown phase Mitman et al. (2023); Cheung et al. (2023).

### v3

@@ V Discussion
In this work we provide straightforward expressions for the first three moments of the news (that is, the displacement, drift, and ballistic memories) in terms of charges and radiative and non-radiative fluxes as well as the first numerical calculation of these terms in the context of a binary black hole merger. Besides providing a hierarchy for these terms in terms of how much they contribute to the various memory effects, we also explore how much the morphology of these contributions match “derivatives of step functions”, which is what they should resemble if one assumes that they predominantly source non-oscillatory memory effects. In particular, we find that while the $\mathcal{F}_{0}$ and $\mathcal{F}_{1}$ flux contributions to the displacement and spin memory effects resemble step- and delta-like functions, the other fluxes break from this expected behavior. A possible explanation for this, however, could simply be that to make these terms exhibit this behavior, then one must perform a correction like the Wald-Zoupas correction that is used for the first moment of the news Wald and Zoupas (2000); Flanagan and Nichols (2017); Grant et al. (2022) or remove the nontrivial QNM contributions that are sourced during the ringdown phase Mitman et al. (2023); Cheung et al. (2023).